i.                   The price elasticity of demand of colour TVs is estimated to be - 2.5 if the prices of a colour TVs is reduced by 20 % how much percentage increase in the quantity of colour TVs sold do you expect?


1
Expert's answer
2021-02-15T17:49:47-0500

By the definition of the price elasticity of demand, we have:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P},""\%\Delta Q=E_d\cdot\%\Delta P=-2.5\cdot(-0.2)\times100\%=50\%."


Therefore, 20% reduction in price will cause 50% increase in the quantity of colour TVs sold.